can some one help me solve this problem? A chemistry experiment requires 60 mL of a 40% alcohol solution. the lab has an 80% alcohol solution and a 30% alcohol solution. how many mL of the two solutions should be mixed to obtain 60 mL of a 40% alcohol solution?
Answer by stanbon(75887)

A chemistry experiment requires 60 mL of a 40% alcohol solution. the lab has an 80% alcohol solution and a 30% alcohol solution. how many mL of the two solutions should be mixed to obtain 60 mL of a 40% alcohol solution?
----------------------
Equation:
alcohol + alcohol = alcohol
0.80x + 0.30(60-x) = 0.40*60
------
80x + 30*60 - 30x = 40*60
50x = 10*60
x = (1/5)*60
x = 12 mL (amt. of 80% solution needed)
----
60-x = 48 mL (amt. of 30% solution needed)
================================================
Cheers,
Stan H.
==============